# Cluster-Enhanced Federated Graph Neural Network for Recommendation

### v1

@@
Abstract
Personal interaction data can be effectively modeled as individual graphs for each user in recommender systems. Graph Neural Networks (GNNs)-based recommendation techniques have become extremely popular since they can capture high-order collaborative signals between users and items by aggregating the individual graph into a global interactive graph. However, this centralized approach inherently poses a threat to user privacy and security. Recently, federated GNN-based recommendation techniques have emerged as a promising solution to mitigate privacy concerns. Nevertheless, current implementations either limit on-device training to an unaccompanied individual graphs or necessitate reliance on an extra third-party server to touch other individual graphs, which also increases the risk of privacy leakage. To address this challenge, we propose a Cluster-enhanced Federated Graph Neural Network framework for Recommendation, named CFedGR, which introduces high-order collaborative signals to augment individual graphs in a privacy preserving manner. Specifically, the server clusters the pretrained user representations to identify high-order collaborative signals. In addition, two efficient strategies are devised to reduce communication between devices and the server. Extensive experiments on three benchmark datasets validate the effectiveness of our proposed methods.
Index Terms:
Federated Learning, Graph Neural Network, Recommender System, Privacy-Preserving, Cluster
@@ I Introduction
In recommendation systems, the interaction data for each user can be effectively represented as a individual graph structures [1]. These personalized graphs capture the unique relationships and interactions between the user and various items, allowing for a more nuanced and accurate modeling of user preferences[2]. The advanced capability of Graph Neural Networks (GNNs) to aggregate individual user interaction graphs into a comprehensive global interactive graph has led to their widespread application in recommender systems [3]. However, this centralized learning methodology necessitates the aggregation and centralized storage of substantial volumes of users’ confidential data[4]. Despite its advantages, this method has several limitations: 1)The data involved is highly sensitive and the transmission of users’ private interaction data beyond local devices poses a significant risk of privacy leakage[5]. 2)Centralized data storage necessitates that third-party devices possess sufficient storage capacity. 3)Stringent data protection regulations, such as the General Data Protection Regulation (GDPR)[6] have been enacted, prohibiting commercial entities from collecting, exchanging, or selling user data without explicit user consent. These limitations make it difficult to deploy centralized learning methods in real-life scenarios.
Fortunately, the Federated Learning (FL) framework proposed by Google[7] addresses these issues. FL allows multiple users to collaboratively train a global model, with users needing to upload only model parameters instead of raw data. In such scenarios, users are confined to constructing first-order subgraphs from their own interaction data, precluding the exploitation of information from other users with similar interests, specifically higher-order collaborative signals. This constraint prevents GNNs from capturing high-order information, leading to degraded recommendation performance[8]. To address this issue, algorithms like FedGNN[9] have been proposed, but it brought about risks of privacy leakage and increased communication burdens. Acquiring advanced collaborative signals while maintaining user privacy has emerged as a critical challenge.
@@ II-A Graph Neural Network
GNN have garnered extensive research attention in recent years owing to their exceptional capability to efficiently model graph data, which has prompted their wide adoption in different graph-based tasks[10], such as node classification[11], link prediction[12], and graph classification[13]. By employing the technique of message passing defined as follows:
@@
where $\mathbf{h}_{v}^{(k+1)}$ represents the feature vector of node $v$ at layer $k+1$, and $\mathbf{h}_{u}^{(k)}$ is the feature vector of node $u$ at layer $k$. The set $\mathcal{N}(v)$ consists of the neighbor nodes of node $v$. The AGGREGATE function collects information from the neighbors, utilizing methods such as summation, averaging, or taking the maximum. $\mathbf{W}$ is a matrix of learnable weights and $\sigma$ signifies the activation function. GNN can iteratively combine the features of neighboring nodes with its own features to update the node’s representation[14]. For example, Kipf et al.[15] proposed a scalable approach for semi-supervised learning on graph-structured data using Graph Convolutional Networks (GCNs). It employs a localized first-order approximation of spectral graph convolutions to efficiently capture local graph structure and node features. Hamilton et al.[16] introduced GraphSAGE, which enhances GNN by sampling and aggregating features from a node’s local neighborhood. Veličković et al. [17] presented Graph Attention Networks (GAT) that employ an attention mechanism to assign different importances to nodes in a neighborhood, allowing the model to focus on the most relevant parts of the graph. He et al.[18] proposed LightGCN, by eliminating feature transformation and nonlinear activation, they simplify traditional GCNs, creating a more efficient model that focuses solely on neighborhood aggregation.
@@ II-B Federated Recommendation Systems
Federated Learning (FL) allows multiple clients to collaboratively train machine learning models while safeguarding data privacy and has seen a significant rise in application to recommendation systems in recent years, which is known as Federated Recommender Systems (FedRecs) [19]. In this paradigm, rather than transmitting their private data, clients engage in the training of local model using data stored locally and subsequently transmit their model parameters or gradients to a central server, which then aggregates this information to derive a comprehensive global model, effectively mitigating the risk of data leakage. FedRecs can generally be categorized into two approaches: federated recommendation based on matrix factorization and federated recommendation based on graph neural networks. Matrix factorization is a technique commonly used in recommendation systems. It works by decomposing the user-item rating matrix into two lower-dimensional matrices, the user matrix U and the item matrix V, such that their product approximates the original rating matrix. For example, Federated Collaborative Filtering (FCF) [20] suggested each client utilizes local data and item embeddings retrieved from the server to update user embeddings, ensuring that user data never leaves the client device. Du,Yongjie et al.[21] introduced a randomized response mechanism to further enhance privacy protection. Recommendation methods based on graph neural networks leverage these networks to derive latent representations of users and items, which are consequently employed for downstream recommendation tasks. For instance, FedSoG[22] adopts a relational attention mechanism to distinguish the contributions of neighboring users and interacted items to the user’s embedding, effectively addressing heterogeneity in the process. A similar work to ours is FedGNN [9]. It utilizes a trusted third-party server and homomorphic encryption to identify users’ neighbors and augment the local subgraph. However, this approach still carries the risk of leaking user privacy and requires higher communication overhead.
@@ III-C Local User-Item Graph Augmentation via Clustering
Firstly, the server that holds all user embeddings performs clustering on these embeddings. To prevent performance degradation caused by inaccurate user embeddings at the initial stage, we conduct clustering after pre-training for $T$ rounds. Numerous clustering algorithms available for this purpose, including the K-means algorithm[23], hierarchical clustering[24], and the DBSCAN algorithm[25]. In our method, we employ the K-means clustering algorithm. After that, users with similar interests are grouped into the same cluster. Except for the user themselves, the remaining users in the cluster are considered the user’s collaborative neighbors. Proceeding with this, the IDs of other users in the cluster are sent to each user by server, and edges are added between the user and the received neighbor nodes. Throughout this process, the server does not receive any private data from the users, nor is there any data sharing between users, which ensures that the local first-order user-item graph can be expanded while maintaining user privacy.
@@ IV-A Experimental Settings and Comparisons
Each dataset is split into training, validation, and test sets with a ratio of $70\%$, $10\%$, and $20\%$, respectively. The learning rate is optimized within the range of {0.01, 0.05, 0.1}, while the decay rate is explored within {0.0001, 0.0005, 0.001}. Each iteration engages 256 users, with embeddings mapped into 128 dimensions. We establish a gradient clipping threshold of 0.2 and set the noise scale to 0.1. Additionally, the $k$-value is set to 200 and 1000 negative samples are utilized. A two-layer Graph Attention Network (GAT) serves as the local GNN. Performance is assessed using two classic metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE), with lower values indicating better results. Baseline models from both centralized and federated learning frameworks are employed to evaluate recommendation efficacy.
@@
IV-B Analysis of Experimental Results
  1. 1.
    TABLE III: Number of neighbors required by FedGNN
    Datasets of neighbors
    Filmtrust 1094
    Douban 815
    MovieLens-100K 860
  2. 2.
  3. 3.
    Despite FedGNN showing marginally better than our model, the difference is minimal, with a maximum improvement of only 0.3%. However, FedGNN necessitates an additional server, which not only heightens the risk of privacy leakage but also increases server deployment costs and communication burdens. Furthermore, as shown in Table III, the number of neighbors that FedGNN needs to transmit is significantly higher than in our model. Our method reduces neighbor communication by approximately 4 to 5 times, achieving a superior balance between communication overhead and recommendation performance.
IV-C Hyperparameter Analysis
TABLE IV: Performance comparison on different values of N
@@ IV-C Hyperparameter Analysis
In this section, we discuss the impacts of hyperparameters, specifically the number of clusters $N$ and the value of top $K$. We conducted experiments on two relatively large datasets, MovieLens-100K and Douban. Table IV demonstrates that performance remains stable across various cluster numbers when the top $K$ is fixed at 40, which alleviates the burden of parameter tuning. Furthermore, as illustrated in Fig. 3, RMSE gradually decreases with increasing $K$. However, once $K$ reaches 300, the reduction in error becomes negligible. To minimize communication overhead, we recommend setting the $K$-value to 200.

### v2

@@
Abstract: Personal interaction data can be effectively modeled as individual graphs for each user in recommender systems. Graph Neural Networks (GNNs)-based recommendation techniques have become extremely popular since they can capture high-order collaborative signals between users and items by aggregating the individual graph into a global interactive graph. However, this centralized approach inherently poses a threat to user privacy and security. Recently, federated GNN-based recommendation techniques have emerged as a promising solution to mitigate privacy concerns. Nevertheless, current implementations either limit on-device training to an unaccompanied individual graphs or necessitate reliance on an extra third-party server to touch other individual graphs, which also increases the risk of privacy leakage. To address this challenge, we propose a Cluster-enhanced Federated Graph Neural Network framework for Recommendation, named CFedGR, which introduces high-order collaborative signals to augment individual graphs in a privacy preserving manner. Specifically, the server clusters the pretrained user representations to identify high-order collaborative signals. In addition, two efficient strategies are devised to reduce communication between devices and the server. Extensive experiments on three benchmark datasets validate the effectiveness of our proposed methods.
Keywords: Federated Learning, Graph Neural Network, Recommender System, Privacy-Preserving, Cluster
@@ I Introduction
In recommendation systems, the interaction data for each user can be effectively represented as a individual graph structures [1, 2, 3]. These personalized graphs capture the unique relationships and interactions between the user and various items, allowing for a more nuanced and accurate modeling of user preferences[4, 5, 6]. The advanced capability of Graph Neural Networks (GNNs) to aggregate individual user interaction graphs into a comprehensive global interactive graph has led to their widespread application in recommender systems [7, 8, 9]. However, the centralized learning methodology necessitates the aggregation and centralized storage of substantial volumes of users’ confidential data[10, 11, 12]. Despite its advantages, this method has several limitations: 1)The data involved is highly sensitive and the transmission of users’ private interaction data beyond local devices poses a significant risk of privacy leakage[13, 14]. 2)Centralized data storage necessitates that third-party devices possess sufficient storage capacity[15]. 3)Stringent data protection regulations, such as the General Data Protection Regulation (GDPR)[16, 17] have been enacted, prohibiting commercial entities from collecting, exchanging, or selling user data without explicit user consent. These limitations make it difficult to deploy centralized learning methods in real-life scenarios.
Fortunately, the Federated Learning (FL) framework proposed by Google[18, 19] addresses these issues. FL allows multiple users to collaboratively train a global model, with users needing to upload only model parameters instead of raw data. In such scenarios, users are confined to constructing first-order subgraphs from their own interaction data, precluding the exploitation of information from other users with similar interests, specifically higher-order collaborative signals[20]. This constraint prevents GNNs from capturing high-order information, leading to degraded recommendation performance[21, 22, 23]. To address this issue, algorithms like FedGNN[24, 25] have been proposed, but it brought about risks of privacy leakage and increased communication burdens. Acquiring advanced collaborative signals while maintaining user privacy has emerged as a critical challenge.
@@ II-A Graph Neural Network
GNN have garnered extensive research attention in recent years owing to their exceptional capability to efficiently model graph data, which has prompted their wide adoption in different graph-based tasks[26, 27], such as node classification[28, 29, 30], link prediction[31, 32], and graph classification[33, 34, 35]. By employing the technique of message passing defined as follows:
@@
where $\mathbf{h}_{v}^{(k+1)}$ represents the feature vector of node $v$ at layer $k+1$, and $\mathbf{h}_{u}^{(k)}$ is the feature vector of node $u$ at layer $k$. The set $\mathcal{N}(v)$ consists of the neighbor nodes of node $v$. The AGGREGATE function collects information from the neighbors, utilizing methods such as summation, averaging, or taking the maximum. $\mathbf{W}$ is a matrix of learnable weights and $\sigma$ signifies the activation function. GNN can iteratively combine the features of neighboring nodes with its own features to update the node’s representation[36, 37, 38]. For example, Kipf et al.[39] proposed a scalable approach for semi-supervised learning on graph-structured data using Graph Convolutional Networks (GCNs). It employs a localized first-order approximation of spectral graph convolutions to efficiently capture local graph structure and node features. Hamilton et al.[40] introduced GraphSAGE, which enhances GNN by sampling and aggregating features from a node’s local neighborhood. Veličković et al. [41] presented Graph Attention Networks (GAT) that employ an attention mechanism to assign different importances to nodes in a neighborhood, allowing the model to focus on the most relevant parts of the graph. He et al.[42] proposed LightGCN, by eliminating feature transformation and nonlinear activation, they simplify traditional GCNs, creating a more efficient model that focuses solely on neighborhood aggregation.
@@ II-B Federated Recommendation Systems
Federated Learning (FL) allows multiple clients to collaboratively train machine learning models while safeguarding data privacy and has seen a significant rise in application to recommendation systems in recent years, which is known as Federated Recommender Systems (FedRecs) [43, 12, 44]. In this paradigm, rather than transmitting their private data, clients engage in the training of local model using data stored locally and subsequently transmit their model parameters or gradients to a central server, which then aggregates this information to derive a comprehensive global model, effectively mitigating the risk of data leakage[45]. FedRecs can generally be categorized into two approaches: federated recommendation based on matrix factorization and federated recommendation based on graph neural networks. Matrix factorization is a technique commonly used in recommendation systems. It works by decomposing the user-item rating matrix into two lower-dimensional matrices[46, 47], the user matrix U and the item matrix V, such that their product approximates the original rating matrix. For example, Federated Collaborative Filtering (FCF) [48] suggested each client utilizes local data and item embeddings retrieved from the server to update user embeddings, ensuring that user data never leaves the client device. Du,Yongjie et al.[49] introduced a randomized response mechanism to further enhance privacy protection. Recommendation methods based on graph neural networks leverage these networks to derive latent representations of users and items, which are consequently employed for downstream recommendation tasks. For instance, FedSoG[50] adopts a relational attention mechanism to distinguish the contributions of neighboring users and interacted items to the user’s embedding, effectively addressing heterogeneity in the process. A similar work to ours is FedGNN [24]. It utilizes a trusted third-party server and homomorphic encryption to identify users’ neighbors and augment the local subgraph. However, this approach still carries the risk of leaking user privacy and requires higher communication overhead.
@@ III-C Local User-Item Graph Augmentation via Clustering
Firstly, the server that holds all user embeddings performs clustering on these embeddings. To prevent performance degradation caused by inaccurate user embeddings at the initial stage, we conduct clustering after pre-training for $T$ rounds. Numerous clustering algorithms available for this purpose, including the K-means algorithm[51, 52], hierarchical clustering[53, 54, 54], and the DBSCAN algorithm[55, 56]. In our method, we employ the K-means clustering algorithm. After that, users with similar interests are grouped into the same cluster. Except for the user themselves, the remaining users in the cluster are considered the user’s collaborative neighbors. Proceeding with this, the IDs of other users in the cluster are sent to each user by server, and edges are added between the user and the received neighbor nodes. Throughout this process, the server does not receive any private data from the users, nor is there any data sharing between users, which ensures that the local first-order user-item graph can be expanded while maintaining user privacy.
@@ IV-A Experimental Settings and Comparisons
Each dataset is split into training, validation, and test sets with a ratio of $70\%$, $10\%$, and $20\%$, respectively. The learning rate is optimized within the range of {0.01, 0.05, 0.1}, while the decay rate is explored within {0.0001, 0.0005, 0.001}. Each iteration engages 256 users, with embeddings mapped into 128 dimensions. We establish a gradient clipping threshold of 0.2 and set the noise scale to 0.1. Additionally, the $k$-value is set to 200 and 1000 negative samples are utilized. A two-layer Graph Attention Network (GAT) serves as the local GNN. Performance is assessed using two classic metrics: Mean Absolute Error (MAE)[5] and Root Mean Square Error (RMSE)[57, 58], with lower values indicating better results. Baseline models from both centralized and federated learning frameworks are employed to evaluate recommendation efficacy.
@@
TABLE III: Performance comparison on different values of N
@@
IV-B Analysis of Experimental Results
TABLE IV: Number of neighbors required by FedGNN
Datasets of neighbors
Filmtrust 1094
Douban 815
MovieLens-100K 860
@@ IV-B Analysis of Experimental Results
  1. 1.
  2. 2.
  3. 3.
    Despite FedGNN showing marginally better than our model, the difference is minimal, with a maximum improvement of only 0.3%. However, FedGNN necessitates an additional server, which not only heightens the risk of privacy leakage but also increases server deployment costs and communication burdens. Furthermore, as shown in Table IV, the number of neighbors that FedGNN needs to transmit is significantly higher than in our model. Our method reduces neighbor communication by approximately 4 to 5 times, achieving a superior balance between communication overhead and recommendation performance.
IV-C Hyperparameter Analysis
In this section, we discuss the impacts of hyperparameters, specifically the number of clusters $N$ and the value of top $K$. We conducted experiments on two relatively large datasets, MovieLens-100K and Douban. Table III demonstrates that performance remains stable across various cluster numbers when the top $K$ is fixed at 40, which alleviates the burden of parameter tuning. Furthermore, as illustrated in Fig. 3, RMSE gradually decreases with increasing $K$. However, once $K$ reaches 300, the reduction in error becomes negligible. To minimize communication overhead, we recommend setting the $K$-value to 200.